OpenStudy (anonymous):

Use the laws of exponents to simplify 5^3/4 * 5^1/8

OpenStudy (anonymous):

i hate adding fractions

myininaya (myininaya):

lol

myininaya (myininaya):

\[5^{\frac{3}{4}+\frac{1}{8}}\]

OpenStudy (anonymous):

...

OpenStudy (anonymous):

this is a multiplication not an addition

myininaya (myininaya):

to multiply numbers with same bases you add the exponents \[5^\frac{6+1}{8}=5^\frac{7}{8}=\sqrt[8]{5^7}\]